Two particles are released from the same height at an interval of 1s. How long after the first particle begins to fall will be the two particles be 10m apart? 1. 1.5s 2. 2s 3. 1.25s 4. 2.5s

Grade:12

1 Answers

John Sharon Sandeep
6 years ago
Let the objects be O1 and O2,when the given situation appears, Let S1=x , therefore S2=x-10. Let t2=t, since O1 started 1s earlier,therefore t1=t+1.☺ Now since S=ut+at^2/2 => S=at^2/2 (as u=o).➡eq1: x=9.8×(t+1)^2 /2 & eq2: x-10=9.8×t2 /2 . Now eq1-eq2 => +10=9.8(t2+1+2t)/2 - 9.8×t2 /2 => 10=9.8(t2+1+2t-t2) /2 => 20=9.8+19.6t => 10.2/19.6=t => t=0.5(approx.)😆 therefore time take for the process= t+1 =1+0.5=1.5s✒